# Tail Distribution of Regret in Optimistic Reinforcement Learning

### v2

@@ Abstract
We derive instance-dependent tail bounds for the regret of optimism-based reinforcement learning in finite-horizon tabular Markov decision processes with unknown transition dynamics. Focusing on a UCBVI-type algorithm, we characterize the tail distribution of the cumulative regret $R_{K}$ over $K$ episodes, rather than only its expectation or a single high-probability quantile. We analyze two natural exploration-bonus schedules: (i) a $K$-dependent scheme that explicitly incorporates the total number of episodes $K$, and (ii) a $K$-independent scheme that depends only on the current episode index. For both settings, we obtain an upper bound on $\mathbb{P}(R_{K}\geq x)$ that exhibits a distinctive two-regime structure: a sub-Gaussian tail starting from an instance-dependent scale $m_{K}^{\mathrm{mb}}$ up to a transition threshold, followed by a sub-Weibull tail beyond that point. We further derive corresponding instance-dependent bounds on the expected regret $\mathbb{E}[R_{K}]$. The proposed algorithm depends on a tuning parameter $\alpha$, which balances the expected regret and the range over which the regret exhibits a sub-Gaussian tail. To the best of our knowledge, our results provide one of the first comprehensive tail-regret guarantees for a standard optimistic algorithm in episodic reinforcement learning.
@@ 1 Introduction
In this paper, we take a step towards closing this gap by studying the concentration of regret for a UCBVI-type algorithm in finite-horizon MDPs. We show that the tail of the cumulative regret, $\mathbb{P}(R_{K}\geq x)$, admits an explicit, non-asymptotic sub-Gaussian or sub-Weibull bound, depending on the value of $x$ and the instance-dependent $Q^{\star}$-gap parameter.
Our contributions are as follows:
@@
  • For the $K$-dependent scheme, we show that the regret exhibits a sub-Gaussian tail on an intermediate regime of $x$ and a sub-Weibull tail beyond $x>\tilde{\mathcal{O}}\!\big(H^{5/2}K^{(1+\alpha)/2}\big)$, where $\alpha\in[0,1]$ is a tuning parameter.
  • For the $K$-independent scheme, we obtain a similar two-regime behavior, where the transition to sub-Weibull happens at $\tilde{\mathcal{O}}\!\big(H^{(5-\alpha)/(2-\alpha)}K^{1/(2-\alpha)}\big)$.
  • We further show that the expected regret is $\mathbb{E}[R_{K}]\;\leq\;\tilde{\mathcal{O}}\!\Big(H^{6}SA(S+K^{\alpha})/\mathrm{gap}^{*}\Big)$, where $\mathrm{gap}^{*}>0$ is the global $Q^{\star}$-gap. The tuning parameter $\alpha$ balances the expected regret and the range over which the regret exhibits a sub-Gaussian tail.
Our results provide the first characterization of the tail of the regret for a standard optimistic algorithm in tabular episodic RL, and they suggest that the entire distribution of the regret—not just its expectation or a single high-probability quantile—can be controlled in an instance-dependent manner determined by the $Q^{\star}$-gap.
@@ 2 Related Work
In the episodic finite-horizon setting considered in this paper, the optimal worst-case scaling was established by the UCBVI algorithm of Azar et al. (2017), which attains regret $\widetilde{\mathcal{O}}\!\big(H\sqrt{SAK}+H\sqrt{KH}\big)$ and derive information-theoretic lower bounds of $\widetilde{\mathcal{O}}\!\big(H\sqrt{SAK}\big)$. A number of works have since refined or extended this result: Efroni et al. (2019) proposed the EULER algorithm, which yields worst-case bounds with improved dependence on the horizon; see also the recent minimax analyses and extensions to factored MDPs, e.g., Tian et al. (2020). These results, however, are primarily worst-case and focus on bounding either the expectation or a single high-probability quantile of the regret, without characterizing the entire tail behavior.
@@
Under the linear MDP assumption with function approximation, He et al. (2021) establish instance-dependent logarithmic regret for LSVI-UCB and UCRL-VTR in linear and linear-mixture MDPs. Velegkas et al. (2022) further refine this picture by designing algorithms that incur only a bounded number of policy switches. Most recently, Zhao et al. (2025) extend these guarantees to adversarial environments.
@@ 3 Problem Formulation
We write $\mathbf{\Pi}^{*}$ for the set of optimal deterministic Markov policies. Our analysis is instance-dependent and exploits a global $Q^{*}$-gap:
$\mathrm{gap}_{h}(s,a)\;$ $\coloneqq\;\Big(Q_{h}^{*}(s,\pi_{h}^{*}(s))-Q_{h}^{*}(s,a)\Big),\qquad\forall(s,h)\in\mathcal{S}\times[H],$
$\mathrm{gap}^{*}\;$ $\coloneqq\;\min_{\begin{subarray}{c}(s,h)\in\mathcal{S}\times[H]\\ a\notin\mathcal{A}_{h}^{*}(s)\end{subarray}}\mathrm{gap}_{h}(s,a),$ (1)
where $\bm{\pi}^{*}\in\mathbf{\Pi}^{*}$ is any optimal Markov policy.
@@
$R_{K}\;=\;\sum_{k=0}^{K-1}\Big(V_{0}^{\bm{\pi}^{*}}(s_{0})-V_{0}^{\bm{\pi}^{k}}(s_{0})\Big),$
where $\bm{\pi}^{*}\in\mathbf{\Pi}^{*}$ is any optimal policy. Our goal is to develop an algorithm that attains low expected cumulative regret and, simultaneously, exhibits a desirable tail distribution for $R_{K}$.
@@ 4 Main Results
In this section, we present our main results. We study the tail behavior of the regret for a UCBVI-type algorithm in the model-based setting. We analyze a UCBVI-type algorithm that uses the following bonus function:
@@
The term $K^{\alpha}$ in the definition of $b_{h}^{k}(\cdot)$ plays the role of a global exploration budget that directly controls the aggressiveness of the bonuses. In the original minimax UCBVI analysis of Azar et al. (2017), as well as in the subsequent instance-dependent analysis of Simchowitz and Jamieson (2019), the exploration terms scale only logarithmically with the time horizon, roughly as $\sqrt{{\log(KSAH)}/{n}}$, which is tailored to control of the expected regret. In contrast, our bonuses scale as $\sqrt{(S\ln 2+\mu K^{\alpha})/n}$ (or $\sqrt{(S\ln 2+\mu(k+1)^{\alpha})/n}$ in the $K$-independent case), so that the uncertainty radius around the empirical model grows polynomially in $K$. This more conservative choice incentivizes sustained exploration even in later episodes: when the algorithm has visited a state–action pair only moderately many times, the bonus remains substantial and pushes the optimistic value estimates upward. From a tail perspective, this is precisely what allows us to derive nontrivial bounds on $\mathbb{P}(R_{K}\geq x)$ for large thresholds $x$, at the price of a controlled polynomial dependence on $K$ inside the bonus which will affect the expected regret.
@@
$\mathbb{P}(R_{K}\geq x)\ \leq\ \exp\!\left(-\frac{\big(x-H\lceil K^{\gamma}\rceil-\,m_{K}^{\mathrm{mb}}\big)_{+}^{\,2}}{50H^{5}K}\right)\;+\;\delta_{K}^{\mathrm{mb}}(\gamma),\qquad\forall x\geq H\lceil K^{\gamma}\rceil,$
@@
$m_{K}^{\mathrm{mb}}\;\coloneqq\frac{32H^{6}SA(0.5S\ln 2+\mu K^{\alpha})}{\mathrm{gap}^{*}},$ (2)
@@
$m_{K}^{\mathrm{mb}}\;\approx\;\mathcal{O}\!\big(H^{6}SA(S+K^{\alpha})/\mathrm{gap}^{*}\big)$
up to logarithmic factors. For thresholds $x$ that are moderately larger than $m_{K}^{\mathrm{mb}}$, the bound $\mathbb{P}(R_{K}\geq x)$ is of sub-Gaussian form with variance proxy $25H^{5}K$, reflecting the Martingale structure of the regret fluctuations after the main exploration phase.
@@
$\mathbb{E}[R_{K}]\ \leq\ H\lceil K^{\gamma}\rceil+m_{K}^{\mathrm{mb}}+6KH^{3}\delta_{K}^{\mathrm{mb}}(\gamma),$
@@
Proposition 1 obtains an instance-dependent upper bound on the expected regret. The first two terms, $m_{K}^{\mathrm{mb}}+H\lceil K^{\gamma}\rceil$, capture the contribution of the baseline instance-dependent level of regret and the initial “burn-in” phase. The final term, proportional to $6KH^{3}\delta_{K}^{\mathrm{mb}}(\gamma)$, accounts for the rare events in which the empirical model falls outside the “good” set at some episodes. In such events the estimated transition kernel $\hat{\mathbb{P}}^{k}$ departs substantially from the true kernel $\mathbb{P}^{*}$ for some $k\in[K]$. Consequently, we cannot ensure that $V^{k}_{0}(s_{0})\geq V^{*}_{0}(s_{0})$ uniformly over $k$, i.e., the optimism property may fail. Thanks to the exponential decay of $\delta_{K}^{\mathrm{mb}}(\gamma)$ in $K^{\alpha}$ (or $K^{\gamma\alpha}$), this residual contribution is negligible: for any fixed $\alpha>0$ and suitable choice of the constant $\mu>0$, the product $K\delta_{K}^{\mathrm{mb}}(\gamma)$ vanishes.
@@
$\mathbb{P}(R_{K}\geq x)\leq\begin{cases}\exp\left(-\tilde{\mathcal{O}}\left(\frac{x^{2}}{H^{5}K}\right)\right)&\text{if }\mathcal{O}(H^{6}SA(S+K^{\alpha})/\mathrm{gap}^{*})\leq x\leq\tilde{\mathcal{O}}(H^{5/2}K^{(1+\alpha)/2})\\ \exp\left(-\tilde{\mathcal{O}}\left(K^{\alpha}\right)\right)&\text{if }x\geq\tilde{\mathcal{O}}(H^{5/2}K^{(1+\alpha)/2})\end{cases}$
@@
$\mathbb{P}(R_{K}\geq x)\leq\begin{cases}\exp\left(-\tilde{\mathcal{O}}\left(\frac{x^{2}}{H^{5}K}\right)\right)&\text{if }\mathcal{O}(H^{6}SA(S+K^{\alpha})/\mathrm{gap}^{*})\leq x\leq\tilde{\mathcal{O}}(H^{(5-\alpha)/(2-\alpha)}K^{1/(2-\alpha)})\\ \exp\left(-\tilde{\mathcal{O}}\left(\frac{x^{\alpha}}{H^{\alpha}}\right)\right)&\text{if }x\geq\tilde{\mathcal{O}}(H^{(5-\alpha)/(2-\alpha)}K^{1/(2-\alpha)})\end{cases}$
Furthermore, $\mathbb{E}[R_{K}]\leq\mathcal{O}(H^{6}SA(S+K^{\alpha})/\mathrm{gap}^{*})$.
@@
The proof follows by setting $\gamma=0$ for KD algorithm, and setting $\gamma=\log_{K}(x/2H)$ for KI algorithm. ∎
@@
Similar instance-dependent rates have been obtained for optimistic $Q$-learning and related model-free methods (Yang et al., 2021; Velegkas et al., 2022). All the algorithms presented in these works depend on the specific choice of confidence level $\delta\in(0,1)$.
@@
$\mathbb{P}(R_{K}\geq x)\;\leq\;\exp\!\Big(-\tilde{\mathcal{O}}\big(x^{2}/(H^{5}K)\big)\Big),$

### v3

@@ Abstract
We derive instance-dependent tail bounds for the regret of optimism-based reinforcement learning in finite-horizon tabular Markov decision processes with unknown transition dynamics. We first study a UCBVI-type (model-based) algorithm and characterize the tail distribution of the cumulative regret $R_{K}$ over $K$ episodes via explicit bounds on $\mathbb{P}(R_{K}\geq x)$, going beyond analyses limited to $\mathbb{E}[R_{K}]$ or a single high-probability quantile. We analyze two natural exploration-bonus schedules for UCBVI: (i) a $K$-dependent scheme that explicitly incorporates the total number of episodes $K$, and (ii) a $K$-independent (anytime) scheme that depends only on the current episode index. We then complement the model-based results with an analysis of optimistic Q-learning (model-free) under a $K$-dependent bonus schedule.
Across both the model-based and model-free settings, we obtain upper bounds on $\mathbb{P}(R_{K}\geq x)$ with a distinctive two-regime structure: a sub-Gaussian tail starting from an instance-dependent scale up to a transition threshold, followed by a sub-Weibull tail beyond that point. We further derive corresponding instance-dependent bounds on the expected regret $\mathbb{E}[R_{K}]$. The proposed algorithms depend on a tuning parameter $\alpha$, which balances the expected regret and the range over which the regret exhibits sub-Gaussian decay.
@@ 1 Introduction
In this paper, we take a step towards closing this gap by studying the concentration of regret for a model-based UCBVI-type algorithm and a model-free $Q$-learning algorithm in finite-horizon MDPs. We show that the tail of the cumulative regret, $\mathbb{P}(R_{K}\geq x)$, admits an explicit, non-asymptotic sub-Gaussian or sub-Weibull bound, depending on the value of $x$ and the instance-dependent $Q^{\star}$-gap parameter. Our contributions are as follows:
@@
  • For the $K$-dependent scheme, we show that the regret exhibits a sub-Gaussian tail on an intermediate regime of $x$ and a sub-Weibull tail beyond $x>\tilde{\mathcal{O}}\!\big(H^{3/2}K^{(1+\alpha)/2}\big)$, where $\alpha\in[0,1]$ is a tuning parameter.
  • For the $K$-independent scheme, we obtain a similar two-regime behavior, where the transition to sub-Weibull happens at $\tilde{\mathcal{O}}\!\big(H^{(3-\alpha)/(2-\alpha)}K^{1/(2-\alpha)}\big)$.
  • We further show that the expected regret is $\mathbb{E}[R_{K}]\;\leq\;\tilde{\mathcal{O}}\!\Big(H^{3}SA(S+K^{\alpha})/\mathrm{gap}^{*}\Big)$, where $\mathrm{gap}^{*}>0$ is the global $Q^{\star}$-gap. The tuning parameter $\alpha$ balances the expected regret and the range over which the regret exhibits a sub-Gaussian tail.
  • Finally, we study the regret of a $Q$-learning-type algorithm in the KD setting, and we establish similar tail and expected regret bounds as above.
Our results provide the first characterization of the tail of the regret for standard optimistic algorithms in tabular episodic RL, and they suggest that the entire distribution of the regret, not just its expectation or a single high-probability quantile, can be controlled in an instance-dependent manner determined by the $Q^{\star}$-gap.
@@ 2 Related Work
In the episodic finite-horizon setting considered in this paper, the optimal worst-case scaling was established by the UCBVI algorithm of Azar et al. (2017), which attains regret $\widetilde{\mathcal{O}}\!\big(H\sqrt{SAK}+H\sqrt{KH}\big)$ and derives information-theoretic lower bounds of $\widetilde{\mathcal{O}}\!\big(H\sqrt{SAK}\big)$. A number of works have since refined or extended this result: Efroni et al. (2019) proposed the EULER algorithm, which yields worst-case bounds with improved dependence on the horizon; see also the recent minimax analyses and extensions to factored MDPs, e.g., Tian et al. (2020). These results, however, are primarily worst-case and focus on bounding either the expectation or a single high-probability quantile of the regret, without characterizing the entire tail behavior.
@@
Under the linear MDP assumption with function approximation, He et al. (2021) establishes instance-dependent logarithmic regret for LSVI-UCB and UCRL-VTR in linear and linear-mixture MDPs. Velegkas et al. (2022) further refine this picture by designing algorithms that incur only a bounded number of policy switches. Most recently, Zhao et al. (2025) extend these guarantees to adversarial environments.
@@ 3 Problem Formulation
Our analysis is instance-dependent and exploits a global $Q^{*}$-gap:
$\mathrm{gap}_{h}(s,a)\;$ $\coloneqq\;\Big(V_{h}^{*}(s)-Q_{h}^{*}(s,a)\Big),\qquad\forall(s,h)\in\mathcal{S}\times[H],$
$\mathrm{gap}^{*}\;$ $\coloneqq\;\min_{\begin{subarray}{c}(s,h)\in\mathcal{S}\times[H]\\ a\notin\mathcal{A}_{h}^{*}(s)\end{subarray}}\mathrm{gap}_{h}(s,a).$ (1)
@@
$R_{K}\;=\;\sum_{k=0}^{K-1}\Big(V_{0}^{*}(s_{0})-V_{0}^{\bm{\pi}^{k}}(s_{0})\Big).$
Our goal is to develop an algorithm that attains low expected cumulative regret and, simultaneously, exhibits a desirable tail distribution for $R_{K}$.
@@ 4 Main Results
In this section, we present our main results. First, in Section 4.1, we study the tail behavior of the regret for a UCBVI-type algorithm in the model-based setting. Next, in Section 4.2, we analyze the regret tail of $Q$-learning as a representative model-free algorithm.
4.1 Model-based tail bound: UCBVI algorithm
We analyze a UCBVI-type algorithm that uses the following bonus function:
@@ 4.1 Model-based tail bound: UCBVI algorithm
The term $K^{\alpha}$ in the definition of $b_{h}^{k}(\cdot)$ plays the role of a global exploration budget that directly controls the aggressiveness of the bonuses. In the original minimax UCBVI analysis of Azar et al. (2017), as well as in the subsequent instance-dependent analysis of Simchowitz and Jamieson (2019), the exploration terms scale only logarithmically with the time horizon, roughly as $\sqrt{{\log(KSAH)}/{n}}$, which is tailored to control the expected regret. In contrast, our bonuses scale as $\sqrt{(S\ln 2+\mu K^{\alpha})/n}$ (or $\sqrt{(S\ln 2+\mu(k+1)^{\alpha})/n}$ in the $K$-independent case), so that the uncertainty radius around the empirical model grows polynomially in $K$. This more conservative choice incentivizes sustained exploration even in later episodes: when the algorithm has visited a state–action pair only moderately many times, the bonus remains substantial and pushes the optimistic value estimates upward. From a tail perspective, this is precisely what allows us to derive nontrivial bounds on $\mathbb{P}(R_{K}\geq x)$ for large thresholds $x$, at the price of a controlled polynomial dependence on $K$ inside the bonus which will affect the expected regret.
@@
$\mathbb{P}(R_{K}\geq x)\ \leq\ \exp\left(-\frac{(x-\lceil K^{\gamma}\rceil H-m_{K}^{\mathrm{mb}})^{2}}{4H(H+1)(2H+1)K/3}\right)\;+\;\delta_{K}^{\mathrm{mb}}(\gamma),\qquad\forall x\geq H\lceil K^{\gamma}\rceil,$
@@
$m_{K}^{\mathrm{mb}}\;\coloneqq\frac{128H^{3}SA(0.5S\ln 2+\mu K^{\alpha})}{\mathrm{gap}^{*}},$ (2)
@@
$m_{K}^{\mathrm{mb}}\;=\;\mathcal{O}\!\big(H^{3}SA(S+K^{\alpha})/\mathrm{gap}^{*}\big).$
For thresholds $x$ that are moderately larger than $m_{K}^{\mathrm{mb}}$, the bound $\mathbb{P}(R_{K}\geq x)$ is of sub-Gaussian form with variance proxy $H(H+1)(2H+1)K/3$, reflecting the martingale structure of the regret fluctuations after the main exploration phase.
@@
$\mathbb{E}[R_{K}]\ \leq\ H\lceil K^{\gamma}\rceil+m_{K}^{\mathrm{mb}}+KH(H+1)\delta_{K}^{\mathrm{mb}}(\gamma),$
@@
Proposition 1 obtains an instance-dependent upper bound on the expected regret. The first two terms, $m_{K}^{\mathrm{mb}}+H\lceil K^{\gamma}\rceil$, capture the contribution of the baseline instance-dependent level of regret and the initial “burn-in” phase. The final term, proportional to $KH(H+1)\delta_{K}^{\mathrm{mb}}(\gamma)$, accounts for the rare events in which the empirical model falls outside the “good” set at some episodes. In such events the estimated transition kernel $\hat{\bm{P}}^{k}$ departs substantially from the true kernel $\bm{P}^{*}$ for some $k\in[K]$. Consequently, we cannot ensure that $V^{k}_{0}(s_{0})\geq V^{*}_{0}(s_{0})$ uniformly over $k$, i.e., the optimism property may fail. Thanks to the exponential decay of $\delta_{K}^{\mathrm{mb}}(\gamma)$ in $K^{\alpha}$ (or $K^{\gamma\alpha}$), this residual contribution is negligible: for any fixed $\alpha>0$ and suitable choice of the constant $\mu>0$, the product $K\delta_{K}^{\mathrm{mb}}(\gamma)$ vanishes.
@@
$\mathbb{P}(R_{K}\geq x)\leq\begin{cases}\exp\left(-\tilde{\mathcal{O}}\left(\frac{x^{2}}{H^{3}K}\right)\right)&\text{if }\mathcal{O}(H^{3}SA(S+K^{\alpha})/\mathrm{gap}^{*})\leq x\leq\tilde{\mathcal{O}}(H^{3/2}K^{(1+\alpha)/2})\\ \exp\left(-\tilde{\mathcal{O}}\left(K^{\alpha}\right)\right)&\text{if }x\geq\tilde{\mathcal{O}}(H^{3/2}K^{(1+\alpha)/2})\end{cases}$
@@
$\mathbb{P}(R_{K}\geq x)\leq\begin{cases}\exp\left(-\tilde{\mathcal{O}}\left(\frac{x^{2}}{H^{3}K}\right)\right)&\text{if }\mathcal{O}(H^{3}SA(S+K^{\alpha})/\mathrm{gap}^{*})\leq x\leq\tilde{\mathcal{O}}(H^{(3-\alpha)/(2-\alpha)}K^{1/(2-\alpha)})\\ \exp\left(-\tilde{\mathcal{O}}\left(\frac{x^{\alpha}}{H^{\alpha}}\right)\right)&\text{if }x\geq\tilde{\mathcal{O}}(H^{(3-\alpha)/(2-\alpha)}K^{1/(2-\alpha)})\end{cases}$
Furthermore, $\mathbb{E}[R_{K}]\leq\mathcal{O}(H^{3}SA(S+K^{\alpha})/\mathrm{gap}^{*})$.
@@
The proof follows by setting $\gamma=0$ for KD algorithm, and setting $\gamma=\log_{K}(x/H)$ for KI algorithm. ∎
@@
In terms of the MDP parameters, this matches the expected regret upper bound presented in Proposition 1. The primary difference is the refined dependence on $K$, which results from a more aggressive exploration strategy. Similar instance-dependent rates have been obtained for optimistic $Q$-learning and related model-free methods (Yang et al., 2021; Velegkas et al., 2022). All the algorithms presented in these works depend on the specific choice of confidence level $\delta\in(0,1)$.
@@
$\mathbb{P}(R_{K}\geq x)\;\leq\;\exp\!\left(-\tilde{\mathcal{O}}\left(\frac{x^{2}}{H^{3}K}\right)\right),$